Am I writing this right? For the written problem. If not can you please correct thank you
The length a rectangle is twice that of its width. If the area of the rectangle is 72 square inches, then find the length and width.
-----
width = w=6
length =X^2
Area = length*width
12=w*L or 2x6=12
72 = 2x*w or 12x36=72
36 = w^2 or 6x6=36
W = 6 inches (width)
2x = 12 inches (length)

Answer by mananth(16946)

width = w=6
length =2x
Area = length*width
12=w*L or 2x6=12
72 = 2x*w or 12x36=72
36 = w^2 or 6x6=36
W = 6 inches (width)
2x = 12 inches (length)